OpenStudy (anonymous):

There are 360 drops in one tablespoon and there are 14.8 mL in one tablespoon. Calculate the number of drops in 1.0 mL.

OpenStudy (anonymous):

24.3 drops (about)

OpenStudy (anonymous):

4. How many molecules of H2O are there in 1.0 g of H2O?

OpenStudy (anonymous):

thats my last question

OpenStudy (anonymous):

uhh, wait 3.34x10^22

OpenStudy (anonymous):

1 mole of H2O= 18 grams 1gram=1/18 mole 1 mole = 6.02x10^23 molecules 6.02x10^23/18=3.34x10^22